Two Questions :
Cody is Twice as old as Evan. Three years ago the sum of their ages was 27. Find the age of each boy.
Jack is twice as old as Lacy. In Three years the sum of their age will be 54. How old are they now.
Your help is greatly appreciated :D

Answer by oberobic(2304)

C = Cody's age
E = Evan's age
C = 2E :: Cody is twice as old as Evan.
Three years ago can be shown as:
C -3 = Cody's age 3 years ago
E -3 = Evan's age 3 years ago
C-3 + E-3 = 27 :: Given
C + E = 33
Substituting C = 2E
2E + E = 33
3E = 33
E = 11
Therefore
C = 2E = 22
How old were they 3 years ago?
C-3 = 19
E-3 = 8
Was the sum of their ages 27 back then?
19 + 8 = 27 ??
Yes.
Answer:
Cody's age is 22.
Evan's age is 11.
.
J = Jack's age
L = Lacy's age
J = 2L
In three years, their ages will be:
J+3 and L+3
Then the sum will be 54.
J+3 + L+3 = 54
J + L = 48
Substitue J = 2L
2L + L = 48
3L = 48
L = 16
J = 2L = 2(16) = 32
How old will they be in 3 years?
J = 35
L = 19
Will the sum of their ages be 54 then?
35 + 19 = 54 ??
Yes.
Answer:
Jack's age is 32 and Lacy's age is 16.
.
Done